PLEASE HELP ASAP:Flying against the jetstream, a jet travels 4240km in 8 hours. Flying with the jetstream, the plane travels 2910km in 3 hours. What is the speed of the jet in still air, and what is the speed of the jetstream?
Thank you so much

Answer by Paul(988)

Let the speed of the wind be y
Let the speed of the plane be y


8x-8y=4240
second equation:
3(x+y)=2910
3x+3y=2910

Multiply the first equation by 3 and the second equatin by 8
24x-24y=12720
24x+24y=23280
-------------
48x=36000
x=750

3(750)+3y=2910
3y=660
y=220

Hence, the speed of the jetstream is 220mph and the speed of the plane is 750mph.
Paul.